Within using a graph students saw that 52% and 96% concentration acetic acid's solutions had the same 1,06 g/cm3 density. Suggest and DESCRIBE, what way we could measure acetic acid's mass fraction in solution which density again is 1,06 g/cm3.
1
Expert's answer
2019-09-20T07:36:25-0400

Mass fraction can be measured by direct titration.

C1*V1=C2*V2

C1 - concentration of alkali solution 1 n

V1 - the amount of alkali spent on titration

V2- acid volume

C2 - acid concentration

С2= С1*V1/V2